# Knowledge Graph Pruning for Recommendation

### v1

@@ Abstract.
Knowledge Graph Pruning, Recommender Systems, Graph Neural Networks
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. INTRODUCTION
Incorporating knowledge graph (KG) as a supplementary resource to support the recommendation system has gained significant popularity. By leveraging KGs, we can obtain a more comprehensive understanding among users, items, and entities, which leads to a better characterization of user preference and item relevance (Wang et al., 2021a, 2019a; Du et al., 2022). KGs can also enhance the transparency and interpretability of recommendation systems, as they provide clear evidence of how we arrive at the recommendation results (Wang et al., 2021a).
To better leverage the KG data, numerous researchers are dedicated to incorporating users, items, and entities into a shared embedding space based on a collaborative knowledge graph (CKG), which aims to manipulate the structured knowledge and produce explainable recommendation results (Wang et al., 2019a). This line of literature adopts the widespread message aggregation scheme by recursively collecting neighborhood information on the CKG to capture the long-range connectivities (Zhou et al., 2020; Wang et al., 2019a; Yang et al., 2023; Wang et al., 2021a). Hence, the valuable knowledge information could be distilled from the KG to boost the collaborative signals over the user-item interactions.
@@
Despite the prevalence, previous literature is based on the steadfast assumption that KGs consistently deliver accurate, reliable, and valuable facts, thereby undoubtedly enhancing recommendation performance (Wang et al., 2021a; Du et al., 2022; Yang et al., 2023). However, in real-world scenarios, large-scale hand-crafted KGs frequently suffer from evident knowledge redundancy and encounter uncontrolled knowledge quality (Xu et al., 2020; Wang et al., 2021b). This not only imposes additional computational overhead but may even adversely undermine recommendation performance. Inspired by this, pruning the KGs is indispensable to better serve recommendation systems with improvements on both efficiency and accuracy.
To this end, we investigate the knowledge graph pruning approach tailored for recommendation systems, which is a never-before-studied problem. Early efforts has been devoted to KG summarization (Doerr and Blenn, 2013; Sacenti et al., 2022; Kang et al., 2022; Brody et al., 2022; Goyal et al., 2020), employing techniques such as cluster-based inference (Doerr and Blenn, 2013; Sacenti et al., 2022) or GNN-based approaches (Kang et al., 2022; Brody et al., 2022; Goyal et al., 2020) to compress the KGs and enhance subsequent utilization efficiency. However, this line of research shares inherent limitations, as the summarized nodes may deviate from the facts that nodes in knowledge graph represent symbolic abstractions of entities from the real-world, thereby impacting the interpretability of the KG. On the contrary, KG pruning aims to obtain a learnable vector to determine the preservation or removal of each triplet, so that we can evaluate their contribution while preserving the original semantics of knowledge. Furthermore, studies in this field solely evaluate the importance of individual attributes or triplets from the structure of KG, without any consideration of recommendation task. Consequently, the pruning result cannot precisely align with the recommendation system, resulting in suboptimal performance.
@@ 2.1. KG-based Recommendation System
Knowledge Graph (KG). Given an entity set $\mathcal{V}$ and a relation set $\mathcal{R}$, a knowledge graph can be denoted as $G=\{(h,r,t)\ \mid h,t\in\mathcal{V},r\in\mathcal{R}\}$. In particular, each triplet $(h,r,t)$ in the KG represents a connection between the head entity $h$ and the tail entity $t$, which is linked by a specific relation $r$. It is noteworthy that $\mathcal{R}$ is comprised of relations from both canonical and inverse directions.
@@
KG-based Recommendation System (KGRS). Given the user set $\mathcal{U}$, the item set $\mathcal{I}$, and their observed interactions $Y$, with the complementary knowledge from $G$, the target of the KGRS task is to learn a function $\mathcal{F}$ capable of predicting the probability that a user $u\in\mathcal{U}$ interacts with a specific item $i\in\mathcal{I}$.
@@ 2.2. Knowledge Graph Pruning for Recommendation
Here we present the definition of knowledge graph pruning for recommendation task (KGPR) which targets at reducing redundant and uninformative edges in knowledge graph $G$ for efficient and effective prediction in KGRS. Specifically, the objective of KGPR task is to learn a binary mask vector $\boldsymbol{s}\in\{0,1\}^{|G|}$ for all triplets, where the $k$-th edge $(h_{k},r_{k},t_{k})$ will be preserved if ${s}_{k}=1$, otherwise removed. To this end, we employ the binary vector $\boldsymbol{s}$ on the original KG to obtain the pruned KG, denoted as $\tilde{G}=G\odot\boldsymbol{s}$. To be specific, the formal definition of KGPR task can be described as:
@@
Knowledge Graph Pruning for Recommendation Systems (KGPR). Given a user set $\mathcal{U}$, an item set $\mathcal{I}$, an interaction matrix $Y$ and a knowledge graph $G$, we aim to learn a binary function $\mathcal{H}$, s.t. $\boldsymbol{s}=\mathcal{H}(\mathcal{U},\mathcal{I},Y,G)$ and $\boldsymbol{s}\in\{0,1\}^{|G|}$, where $\boldsymbol{s}$ is a binary mask vector. Subsequently, we can acquire the pruned KG, denoted as $\tilde{G}=G\odot\boldsymbol{s}$. According to the pruned KG, the objective of KGPR is to maximize the predictive accuracy of the unobserved user-item interactions in KGRS.
@@ 3. METHODOLOGY
In this section, we will explain the KGTrimmer model to achieve knowledge graph pruning for recommendation in detail, which is composed of the following modules, as shown in Figure 2:
@@
  • Pruned Knowledge Graph Generation. We finally aggregate the candidate masks during the training process to obtain a stable and robust pruned knowledge graph.
@@ 3.1. Dual-view Importance Evaluator
In this subsection, we will introduce how we construct the dual-view KGTrimmer to filter out the effective knowledge that is really beneficial to the collaborative filtering. Specifically, we derive the importance score for each entity from two views, i.e. collective view and holistic view. For the collective view, we try to consider the user-item interactions to compute the collective importance scores, which show the specific interacted users’ preference on the entities. For the holistic view, a learnable mask vector is applied which expects to identify the valueless entities from inherent properties or overall popularity, which is known as the holistic importance scores. After that, we balance the influence of the collective and holistic views and aggregate their importance scores to obtain the aggregated importance scores on all KG entities.
@@ 3.1.1. Collective View
Users express personalized preferences over items while the items are linked with the entities. Along this line, we realize that the essence of user interest lies in the entities within the KG. Motivated by the concept of collective intelligence, we take advantage of these abundant user-item interactions and consider them as the results of a social discourse. Coupled with the KG, we collect and aggregate their diverse opinions to extract the consensus and determine the importance of each entity for shaping their user profiles in a collective view. To this end, we intuitively construct a real-value based user-entity matrix to better represent the correlation between the users and entities. After the construction, each entity receives the attention from all related users to generate a score, a.k.a. collective score. We delve into the details as follows:
User-Entity Matrix Construction. For a more streamlined computational process, we gather data from collaborative signals and knowledge graph to assemble the user-entity matrix. To be more specific, we first follow the prior studies (Wang et al., 2019a) to build a collaborative knowledge graph by combining the user-item bipartite graph and the knowledge graph. For simplicity, we denote the matrix as $Q\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{V}|}$, where $|\mathcal{U}|$ and $|\mathcal{V}|$ is the number of users and entities respectively. Then we define :
@@
Due to the knowledge explosion, an entity may reach uncountable users, thereby leading to the substantial computational cost. As a result, we employ a more effective strategy to consider the path from user $u$ to entity $v$. That is to say, if an entity $v$ maintains more than $k$ users, i.e. $|Q_{v}|>k$, we implement a random sampling strategy to set $Q_{uv}=1$ for the $k$ selected users, while setting $Q_{uv}=0$ for the rest. Here $|Q_{v}|$ represents the number of paths that end with entity $v$.
@@
(2) $s_{v}^{c}$ $=\frac{1}{|\mathcal{N}_{v}^{Q}|}\sum_{u\in\mathcal{N}_{v}^{Q}}sim(\boldsymbol{e}_{u},\boldsymbol{e}_{v}^{\prime}),$
$\boldsymbol{e}_{v}^{\prime}$ $=\frac{1}{|\mathcal{N}_{v}^{\mathcal{R}}|}\sum_{r\in\mathcal{N}_{v}^{\mathcal{R}}}\boldsymbol{e}_{r}\odot\boldsymbol{e}_{v},$
where $\mathcal{N}_{v}^{\mathcal{R}}=\{r|(h,r,v)\in G\}$, $sim$ is the similarity function, $\odot$ denotes the element-wise product, $\boldsymbol{e}_{u}$, $\boldsymbol{e}_{v},\boldsymbol{e}_{r}$ are the embeddings of user $u$, entity $v$ and relation $r$, respectively, and $\boldsymbol{e}_{v}^{\prime}$ is the entity embedding which absorbs the information of relations, and $s_{v}^{c}$ is the collective importance score for each entity $v$. Formally, we define the similarity function as :
(3) $sim(\boldsymbol{e}_{u},\boldsymbol{e}_{v}^{\prime})$ $=max(0,\frac{\boldsymbol{e}_{u}^{T}\boldsymbol{e}_{v}^{\prime}}{\|\boldsymbol{e}_{u}\|\|\boldsymbol{e}_{v}^{\prime}\|}),$
We use the modified cosine similarity function for two reasons: (1) We can constrain the output to be within the range of 0 to 1, allowing it to be utilized as a score. (2) Entities with negative similarity are deemed noisy or erroneous nodes, therefore will be assigned the score of zero and prioritized for removal. Therefore, the collective score is endowed with the ability to jointly identify both noisy and valueless nodes. Consequently, we obtain the collective importance scores as:
@@
which significantly demonstrates interacted users’ attention on KG entities based on collaborative signals.

### v2

@@ Abstract.
Knowledge Graph Pruning, Recommendation System, Graph Sparsification
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Recommender systems
@@ 1. INTRODUCTION
Incorporating knowledge graph (KG) as a supplementary resource to support the recommendation system has gained significant popularity. By leveraging KGs, we can obtain a more comprehensive understanding among users, items, and entities, which leads to a better characterization of user preference and item relevance (Wang et al., 2021b, 2019a; Du et al., 2022). KGs can also enhance the transparency and interpretability of recommendation systems, as they provide clear evidence of how we arrive at the recommendation results (Wang et al., 2021b).
To better leverage the KG data, numerous researchers are dedicated to incorporating users, items, and entities into a shared embedding space based on a collaborative knowledge graph (CKG), which aims to manipulate the structured knowledge and produce explainable recommendation results (Wang et al., 2019a). This line of literature adopts the widespread message aggregation scheme by recursively collecting neighborhood information on the CKG to capture the long-range connectivities (Zhou et al., 2020; Wang et al., 2019a; Yang et al., 2023; Wang et al., 2021b). Hence, the valuable knowledge information could be distilled from the KG to boost the collaborative signals over the user-item interactions.
@@
Despite the prevalence, previous literature is based on the steadfast assumption that KGs consistently deliver accurate, reliable, and valuable facts, thereby undoubtedly enhancing recommendation performance (Wang et al., 2021b; Du et al., 2022; Yang et al., 2023). However, in real-world scenarios, large-scale hand-crafted KGs frequently suffer from evident knowledge redundancy and encounter uncontrolled knowledge quality (Xu et al., 2020; Wang et al., 2021c). This not only imposes additional computational overhead but may even adversely undermine recommendation performance. Inspired by this, pruning the KGs is indispensable to better serve recommendation systems with improvements on both efficiency and accuracy.
To this end, we investigate the knowledge graph pruning approach tailored for recommendation systems, which is a never-before-studied problem. Early efforts has been devoted to KG summarization (Doerr and Blenn, 2013; Sacenti et al., 2022; Kang et al., 2022; Brody et al., 2022; Goyal et al., 2020), employing techniques such as cluster-based inference (Doerr and Blenn, 2013; Sacenti et al., 2022) or GNN-based approaches (Kang et al., 2022; Brody et al., 2022; Goyal et al., 2020) to compress the KGs and enhance subsequent utilization efficiency. However, this line of research shares inherent limitations, as the summarized nodes may deviate from the facts that nodes in knowledge graph represent symbolic abstractions of entities from the real-world, thereby impacting the interpretability of the KG. On the contrary, KG pruning aims to obtain a learnable vector to determine the preservation or removal of each triplet, so that we can evaluate their contribution while preserving the original semantics of knowledge. Recently, Graph Sparsification(Li et al., 2022; Zhang et al., 2024a; Yu et al., 2022; Chen et al., 2021) has come into the spotlight, which attempts to remove noisy edges to obtain a robust graph structure. However, studies in this field solely evaluate the importance of individual attributes or triplets from the structure of KG, without any consideration of recommendation task. Consequently, the pruning result cannot precisely align with the recommendation system, resulting in suboptimal performance.
@@
2. RELATED WORK
2.1. Knowledge-Aware Recommendation
Path-based models (Catherine and Cohen, 2016; Hu et al., 2018; Ma et al., 2019; Sun et al., 2018; Wang et al., 2018, 2019b) focus on the patterns of the long-connectivity among users and entities via items. Using RippleNet (Wang et al., 2018) as an illustrative example, this model initiates with users as seeds as disseminates their information in a ripple effect to gather the preferences of all related nodes. Nonetheless, the intricate nature of the graph structure renders RippleNet prone to information explosion. Other methods, such as meta-path (Hu et al., 2018; Jin et al., 2020), are delicately designed to capture the user preference on certain entities, and thus shed some light on the interpretation. However, these schemes are gradually fading out due to the heavy dependency on pattern quality.
Embedding-based models (Ai et al., 2018; Cao et al., 2019; Huang et al., 2018; Wang et al., 2020; Zhang et al., 2016) attempt to incorporate the semantic information in knowledge graph. They equip items with the KGE (knowledge graph embedding) techniques to discover latent relationships among entities. For instance, KTUP (Cao et al., 2019) believes there is an implicit translational relationship between users and items, where the differences between them can be identified as user preferences. To this end, KTUP adopts the TransH (Wang et al., 2014) to resolve the N-to-N issue and establishes a multi-task task to align the information on the side of the knowledge graph. Despite achieving impressive performance on recommendation tasks, these models overlook the higher-order connectivity within the graph.
GNN-based models (Wang et al., 2021b, 2019c; Du et al., 2022; Wei et al., 2022) target at aggregating messages from neighbors to enrich the ego node embedding. which inherits the propagation mechanism of GNN(Zhou et al., 2020). Typically, KGIN (Wang et al., 2021b) proposes an end-to-end fashion to learn the latent user intents and provide more insight on interpretability. KGRec(Yang et al., 2023) proposes a rational-aware masking mechanism to identify the most important information in the KG, which mitigate the influence of the noisy and irrelevant triplets. Although they have achieved stunning success, all of them suffer from extremely high computational costs due to the cumbersome KG. Along this line, we propose a simple yet effective method to prune the undesirable nodes and maintain comparable performance on various models and datasets.
2.2. Knowledge Graph Error Detection
With the growing adoption of LLM-based knowledge graph construction, concerns regarding the quality of these graphs have amplified, leading to an increased emphasis on the development of error detection techniques for graphs. As a result, numerous researchers dedicated significant efforts to studying how to detect erroneous information within knowledge graphs. For example, CKRL(Xie et al., 2018) integrated the local triplets and global path confidence to detect the possible noises. KGTtm(Jia et al., 2019) borrowed the idea from PageRank and proposed resource allocation mechanism to measure the trustworthiness of triplets. SUKE(Wang et al., 2021d) considered the uncertain information and structure in the knowledge graph to predict the rationality of triplets. REFORM(Wang et al., 2021a) leveraged the meta-learning skills to collect the knowledge from meta-tasks and predict the triplets confidence under few-shot settings. CAGED(Zhang et al., 2022) adopted the multi-view contrastive learning technique which is tailored for the knowledge graph error detection.
Although the aforementioned techniques are effective in recognizing errors and noisy triplets within a KG, they fall short in detecting valueless information. Large KGs are inundated with such valueless triplets which offer no benefit to downstream tasks while consuming valuable computational resources. Additionally, these kinds of information would prevent models from identifying truly valuable triplets. In this paper, we are committed to identifying these valueless nodes to provide a higher-quality graph for downstream tasks.
2.3. Graph Summarization
Researchers have devoted themselves to another the to slim the KGs, called Graph Summarization, which compresses the vast KG into a compact one(Liu et al., 2018; Shabani et al., 2023; Doerr and Blenn, 2013; Sacenti et al., 2022; Belth et al., 2020; Kang et al., 2022; Brody et al., 2022; Goyal et al., 2020). One brunch of these approaches leverage machine learning or statistical rules to aggregate nodes. Some works (Doerr and Blenn, 2013; Sacenti et al., 2022) adopt clustering techniques to aggregate neighbors into a super node. Despite the prevalence of these methods, they still are incapable to incorporate the collaborative signals. KGist(Belth et al., 2020) tries to capture the meta-path rules according to the Minimum Description Length principle (MDL). By applying these rules, abnormal edges are identified and eliminated. Another stream of works turns to GNN for help(Kang et al., 2022; Brody et al., 2022; Goyal et al., 2020). For instance, PeGaSus (Kang et al., 2022) optimizes the summarization of large-scale graphs by considering both user preferences and interests, along with the structural properties of the graph. Regardless of remarkable achievements, KG summarization skills still face undesirable limitations. On the one hand, summarized nodes from the aggregated nodes may deviate from the facts that nodes in knowledge graph represent symbolic abstractions of entities from the real-world. On the other hand, they all put emphasis on the role of local neighbors, the signals from the holistic view are greatly neglected. Along this line, we launched an unexplored task knowledge graph pruning for the recommender system and proposed a novel dual-view-based framework to address the two issues above elegantly.
2.4. Knowledge Graph Pruning
Furthermore, the demand for efficiency has sparked the development of Graph Sparsification, which simplifies graph structures to improve computational efficiency. Early work focused on graph partitioning and clustering tasks, which rely on the theoretical analysis of graph properties to guide the sparsification process, such as (Batson et al., 2013; Althöfer et al., 1990). With the rise of GNNs, pruning methods have also emerged. UGS (Chen et al., 2021) seeks lottery tickets(Frankle and Carbin, 2019) in both the graph structure and network parameters to prune task-irrelevant edges and model parameters. Subsequently, AdaGLT(Zhang et al., 2024b) proposed a layer-wise pruning approach that integrates the training and pruning processes, significantly enhancing both efficiency and performance. MoE(Zhang et al., 2024a) further borrowed the concept of Mixture of Experts(Jacobs et al., 1991), using multiple expert models to evaluate the importance of pruned edges. Although these methods have achieved impressive results in graph pruning, they struggle to effectively incorporate collaborative filtering signals, making it challenging to adapt them to recommendation tasks.
@@ 3.1. KG-based Recommendation System
Knowledge Graph (KG). Given an entity set $\mathcal{V}$ and a relation set $\mathcal{R}$, a knowledge graph can be denoted as $\mathcal{G}=\{(h,r,t)\ \mid h,t\in\mathcal{V},r\in\mathcal{R}\}$. In particular, each triplet $(h,r,t)$ in the KG represents a connection between the head entity $h$ and the tail entity $t$, which is linked by a specific relation $r$. It is noteworthy that $\mathcal{R}$ is comprised of relations from both canonical and inverse directions.
@@
KG-based Recommendation System (KGRS). Given the user set $\mathcal{U}$, the item set $\mathcal{I}$, and their observed interactions $Y$, with the complementary knowledge from $\mathcal{G}$, the target of the KGRS task is to learn a function $\mathcal{F}$ capable of ranking the set of items $\mathcal{I}$ that a user $u\in\mathcal{U}$ has the potential to interact with, i.e. Top-K recommendation task.
@@ 3.2. Knowledge Graph Pruning for Recommendation
Here we present the definition of knowledge graph pruning for recommendation task (KGPR) which targets at reducing redundant and uninformative edges in knowledge graph $\mathcal{G}$ for efficient and effective prediction in KGRS. Specifically, the objective of KGPR task is to learn a binary mask vector $\boldsymbol{s}\in\{0,1\}^{|\mathcal{G}|}$ for all triplets, where the $k$-th edge $(h_{k},r_{k},t_{k})$ will be preserved if ${s}_{k}=1$, otherwise removed. To this end, we employ the binary vector $\boldsymbol{s}$ on the original KG to obtain the pruned KG, denoted as $\tilde{\mathcal{G}}=\mathcal{G}\odot\boldsymbol{s}$. To be specific, the formal definition of KGPR task can be described as:
@@
Knowledge Graph Pruning for Recommendation Systems (KGPR). Given a user set $\mathcal{U}$, an item set $\mathcal{I}$, an interaction matrix $Y$ and a knowledge graph $\mathcal{G}$, we aim to learn a binary function $\mathcal{H}$, s.t. $\boldsymbol{s}=\mathcal{H}(\mathcal{U},\mathcal{I},Y,\mathcal{G})$ and $\boldsymbol{s}\in\{0,1\}^{|\mathcal{G}|}$, where $\boldsymbol{s}$ is a binary mask vector. Subsequently, we can acquire the pruned KG, denoted as $\tilde{\mathcal{G}}=\mathcal{G}\odot\boldsymbol{s}$. According to the pruned KG, the objective of KGPR is to maximize the predictive accuracy of the unobserved user-item interactions in KGRS.
@@ 4. METHODOLOGY
In this section, we begin by introducing the formal formulation of the KG pruning task to gain a comprehensive understanding of overall framework. Following this, we introduce the KGTrimmer model to achieve knowledge graph pruning for recommendation in detail.
4.1. Overall Framework
In traditional knowledge graph pruning tasks, researchers often assign a score to each node, which represents its importance. Along this line, we still lay stress on obtaining the final score of each node. In addition, we also aim to maintain the performance of recommendation system after pruning low-informative nodes, i.e. optimize the recommendation system with the pruned knowledge graph. Therefore, there is an urgent need to analyse the task to provide a clearer formula to optimize these two objectives simultaneously. In summary, we formulate the KGPR task as an optimization problem, which is to maximize the probability gap between the positive item $i$ and negative item $j$ for user $u$ given a knowledge graph $\mathcal{G}$, with respect to the estimated mask scores $\hat{\boldsymbol{s}}$:
(1) $P_{\theta}(i|u,\mathcal{G})=P_{\theta}(i,|u,\mathcal{G},\hat{\boldsymbol{s}})P_{\theta}(\hat{\boldsymbol{s}}|u,\mathcal{G})$
where $\theta$ is the parameters of the proposed model, $\hat{\boldsymbol{s}}$ represents the estimated mask scores for all node and $\mathcal{G}$ is the collaborative knowledge graph. Here, $P_{\theta}(i|u,\mathcal{G})$ means the probability of user $u$ interacting with the item $i$ (Rendle et al., 2012). Concretely, The latter term $P_{\theta}(\hat{\boldsymbol{s}}|u,\mathcal{G})$ is to estimate the importance of nodes in knowledge graph $\mathcal{G}$, which can be recognized as importance evaluator. The former term $P_{\theta}(i|u,\mathcal{G},\hat{\boldsymbol{s}})$ is the probability that user $u$ favors the positive item $i$ given the importance score $\hat{\boldsymbol{s}}$ and knowledge graph $\mathcal{G}$, which can be regarded as importance-aware predictor.
It is worth noting that our ultimate goal is to gather the determined score $\hat{\boldsymbol{s}}$ to locate the valueless nodes, making it rational to prune the knowledge graph in the subsequent process. Along this line, this formulation breaks the conventional knowledge-aware recommendation task into two parts and seamlessly borrows the signal from collaborative information to guide the learning process of scores. Furthermore, the desired score set $\boldsymbol{s}\in\{0,1\}^{|\mathcal{G}|}$ is the binary vector, whereas our estimated score set $\hat{\boldsymbol{s}}$ is defined as continuous vector for streamlining the gradient calculation in the training process. To bridge this gap, we design a Pruned Knowledge Graph Generation module to binarize the estimated score which serves as the justification for pruning the knowledge graph. In a nutshell, we elaborate on the details of our framework, which is composed of the following modules, as shown in Figure 2:
@@ 4.1. Overall Framework
  • Pruned Knowledge Graph Generation. We finally aggregate the candidate masks during the training process and binarize the estimated mask vector to obtain a stable and robust pruned knowledge graph.
4.2. Initialization
In line with prior work on KG-aware recommendation (Wang et al., 2021b; Du et al., 2022; Wang et al., 2019a), we assign embedding vector $\boldsymbol{e}^{0}_{u}$, $\boldsymbol{e}^{0}_{v}$ and $\boldsymbol{e}_{r}$ to each user $u\in\mathcal{U}$, entity $v\in\mathcal{V}$ and relation $r\in\mathcal{R}$, respectively. As mentioned above, we omit the item embedding since it is already included in the entity set. Thus, we can construct three matrices:
(2) $\boldsymbol{E}_{u}=\{\boldsymbol{e}^{0}_{u_{1}},\boldsymbol{e}^{0}_{u_{2}}\,\ldots,\boldsymbol{e}^{0}_{u_{|\mathcal{U}|}}\},\quad\boldsymbol{E}_{v}=\{\boldsymbol{e}^{0}_{v_{1}},\boldsymbol{e}^{0}_{v_{2}}\,\ldots,\boldsymbol{e}^{0}_{v_{|\mathcal{V}|}}\},\quad\boldsymbol{E}_{r}=\{\boldsymbol{e}_{r_{1}},\boldsymbol{e}_{r_{2}}\,\ldots,\boldsymbol{e}_{r_{|\mathcal{R}|}}\}$
Here, $|\mathcal{U}|,|\mathcal{V}|,|\mathcal{R}|$ represents the number of the user set $\mathcal{U}$, entity set $\mathcal{V}$ and relation set $\mathcal{R}$, respectively. It is worth noting that, these matrices are initialized via Xavier initializer (Glorot and Bengio, 2010) which guarantees the orthogonality of the initial vectors in the matrix. Taking the user matrix as an example, we have following observations at the initialization step:
(3) $(\boldsymbol{e}^{0}_{u_{j}})^{T}\boldsymbol{e^{0}_{u_{k}}}=0,\quad j\neq k$
where $u_{j}\in\mathcal{U}$ and $u_{k}\in\mathcal{U}$ represent any arbitrary users. Subsequently, we employ one-hot vectors $\boldsymbol{ID}$ for each user, entity and relation, and retrieve their corresponding embeddings:
(4) $\boldsymbol{e}^{0}_{u}=\boldsymbol{E}_{u}^{T}\cdot\boldsymbol{ID}_{u},\quad\boldsymbol{e}^{0}_{v}=\boldsymbol{E}_{v}^{T}\cdot\boldsymbol{ID}_{v},\quad\boldsymbol{e}_{r}=\boldsymbol{E}_{r}^{T}\cdot\boldsymbol{ID}_{r}$
@@ 4.3. Dual-view Importance Evaluator
In this subsection, we attempt figure out the solution to accurately estimating the score $\hat{\boldsymbol{s}}$, i.e. $P_{\theta}(\hat{\boldsymbol{s}}|\mathcal{G})$. Along this line, we will introduce how we construct the dual-view KGTrimmer to filter out the effective knowledge that is really beneficial to the collaborative filtering. To be specific, we derive the importance score $\hat{\boldsymbol{s}}$ for each entity from two views, i.e. collective view and holistic view. For the collective view, we try to consider the user-item interactions to compute the collective importance scores, which show the specific interacted users’ preference on the entities. For the holistic view, a learnable mask vector is applied which expects to identify the valueless entities from inherent properties or overall popularity, which is known as the holistic importance scores. After that, we balance the influence of the collective and holistic views and aggregate their importance scores to obtain the aggregated importance scores on all KG entities.
@@ 4.3.1. Collective View
Users express personalized preferences over items while the items are linked with the entities. Along this line, we realize that the essence of user interest lies in the entities within the KG. Motivated by the concept of collective intelligence, we take advantage of these abundant user-item interactions and consider them as the results of a social discourse. Coupled with the KG, we collect and aggregate their diverse opinions to extract the consensus and determine the importance of each entity for shaping their user profiles in a collective view. To this end, we intuitively construct a real-value based user-entity matrix to better represent the correlation between the users and entities. After the construction, each entity receives the attention from all related users to generate a score, a.k.a. collective score. We delve into the details as follows.
User-Entity Matrix Construction. We first gather data from collaborative signals and knowledge graph to assemble the user-entity matrix. For one thing, it is impractical for all entities being aware of the preference of all users, which would result in unbearably high costs. For another, inspired by Monte Carlo algorithm, we do not need to traverse all users related to the entity $v$ to obtain the $\hat{s}_{v}$. Instead, by adopting a sampling strategy, we can approximate the overall rating of $v$ by examining the scores from the sampled users. Along this line, we follow the prior studies (Wang et al., 2019a) to build a collaborative knowledge graph by combining the user-item bipartite graph and the knowledge graph. For simplicity, we denote the matrix as $Q\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{V}|}$, where $|\mathcal{U}|$ and $|\mathcal{V}|$ is the number of users and entities respectively. Then we define:
@@
Next, we employ a more effective Monte Carlo strategy to sample the path from user $u$ to entity $v$. That is to say, if an entity $v$ maintains more than $k$ users, i.e. $|Q_{v}|>k$, we implement a random sampling strategy to set $Q_{uv}=1$ for the $k$ selected users, while setting $Q_{uv}=0$ for the rest. Here $|Q_{v}|$ represents the number of paths that end with entity $v$.
Based on the Monte Carlo algorithm, the mean derived from numerous random samples closely approximates the true distribution’s mean. Consequently, by sampling $k$ users, we can estimate the overall user attitude towards this entity, with a larger $k$ resulting in a narrower gap between the estimated and true means. Nevertheless, an increase in k also elevates the computational burden. Thus, a reasonable choice of $k$ is made to optimize the trade-off between precision and computational efficiency.
@@
(6) $s_{v}^{c}$ $=\frac{1}{|\mathcal{N}_{v}^{Q}|}\sum_{u\in\mathcal{N}_{v}^{Q}}sim(\boldsymbol{e}_{u}^{0},\boldsymbol{e}_{v}^{\prime}),$
$\boldsymbol{e}_{v}^{\prime}$ $=\frac{1}{|\mathcal{N}_{v}^{\mathcal{R}}|}\sum_{r\in\mathcal{N}_{v}^{\mathcal{R}}}\boldsymbol{e}_{r}\odot\boldsymbol{e}_{v}^{0},$
where $\mathcal{N}_{v}^{\mathcal{R}}=\{r|(h,r,v)\in G\}$, $sim$ is the similarity function, $\odot$ denotes the element-wise product, $\boldsymbol{e}_{u}^{0}$, $\boldsymbol{e}_{v}^{0},\boldsymbol{e}_{r}$ are the embeddings of user $u$, entity $v$ and relation $r$, respectively, and $\boldsymbol{e}_{v}^{\prime}$ is the entity embedding which absorbs the information of relations, and $s_{v}^{c}$ is the collective importance score for each entity $v$. Formally, we define the similarity function as :
(7) $sim(\boldsymbol{e}_{u}^{0},\boldsymbol{e}_{v}^{\prime})$ $=max(0,\frac{(\boldsymbol{e}_{u}^{0})^{T}\boldsymbol{e}_{v}^{\prime}}{\|\boldsymbol{e}_{u}^{0}\|\|\boldsymbol{e}_{v}^{\prime}\|}),$
We use the modified cosine similarity function for the following considerations: (1) This method naturally limits the output to a range between 0 and 1 and can be considered as a form of dot product followed by normalization. (2) This similarity calculation method ensures that the initial values of the collective score $s^{c}_{v}$ are around zero, owing to the fact that the initialization of each matrix is initialized independently. This guarantees that our scores all start from the same point, unlike other normalization methods, such as min-max normalization, which can introduce biases in the scores at the initial stage. Such biases could significantly impact the results of subsequent training. (3) Entities with negative similarity are deemed noisy or erroneous nodes, therefore will be assigned the score of zero and prioritized for removal. Therefore, the collective score is endowed with the ability to jointly identify both noisy and valueless nodes. Consequently, we obtain the collective importance scores as:
@@
which demonstrates interacted users’ attention on KG entities based on collaborative signals.